# GP+: A Python Library for Kernel-based learning via Gaussian Processes

### v1

@@ 4.3 Gaussian Processes with Mixed Basis Functions
The parametric mean and covariance functions in Equation 3b can be formulated in many ways. In this regard, most advancements have focused on designing the kernel (e.g., the ones we develop in Equations 13b and 31 or deep kernels [91, 92]) since it significantly affects the performance of the resulting GP model. However, the mean function in Equation 3a plays an important role in many applications that involve, e.g., extrapolation [93], fusing multi-source data, or identifying model form errors.
Existing techniques typically leverage polynomials (in combination with other analytic functions such as $\sin(\cdot),\log(\cdot),\cdots$) or neural networks in designing $m(\boldsymbol{x};\boldsymbol{\beta})$. In GP+, we extend these methods to seamlessly include the categorical variables in the mean function. Specifically, our idea is to feed the learnt representations of $\boldsymbol{t}$ and $s$ into the mean function instead of the original categorical variables, i.e., we reparameterize $m(\boldsymbol{x},\boldsymbol{t},s;\boldsymbol{\beta})$ to $m(\boldsymbol{x},\boldsymbol{h},\boldsymbol{z};\boldsymbol{\beta})$, see Figure 5. A major difference between our reparameterization and other alternatives (such as an NN whose inputs are one-hot encoded representation of $\boldsymbol{t}$ and $s$) is that our mean and covariance functions are directly coupled since the latent variables used in $m(\boldsymbol{x},\boldsymbol{h},\boldsymbol{z};\boldsymbol{\beta})$ are parameterized in the kernel, i.e., $\boldsymbol{h}=f_{h}(\boldsymbol{\pi}_{t};\boldsymbol{\theta}_{h})$ and $\boldsymbol{z}=f_{z}(\boldsymbol{\pi}_{s};\boldsymbol{\theta}_{z})$.
Figure 5: Multi-fidelity modeling via mixed basis functions: Two generic options are defined in GP+ for building the mixed bases: $(1)$ predetermined bases where multiple bases like polynomial, $sin(\cdot)$ and $cos(\cdot)$ can be defined for each data source, $(2)$ FFNNs with user-defined architectures. All the parameters of the mean functions have a normal prior and are jointly learned through MAP.
Figure 5 illustrates the two generic options that we define in GP+ for building mixed bases. The first option builds the mean function based on pre-determined bases that can include polynomials, $sin(\cdot),log(\cdot)$ or any other analytic functions. The second option is based on a fully connected FFNN whose architecture (e.g., number of hidden layers and their sizes) should be designed by the user. The size of the input layer of the NN depends on the dimensionality of $\boldsymbol{x},\boldsymbol{h},$ and $\boldsymbol{z}$ while its output layer size depends on the response dimensionality (hence the output size is $1$ for a single-response dataset).
@@
The results are illustrated in Figure 6 and indicate that the second GP emulates both sources better than the first GP in both interpolation and extrapolation. As it can be seen in Table 4 the true model form error is $10.3x^{2}-0.7x+$ 1 while the discovered one with the second GP is $0.99390.2981x^{2}-0.7059x+$ 0.9939. We attribute the small differences between these two functions primarily to the fact that the training data is very small and noisy. As also shown in Figure 6 we observe that the inclusion of the mixed bases affects the learnt encoding for $s$ whose two levels are mapped to distant latent points in the first GP but close-by points in the second GP. This behavior indicates that the entire model form error in the second GP can be obtained by comparing the mean functions associated with the two sources.
@@ 4.4 Inverse Parameter Learning for Computer Models
Most computer models are built to be applicable to a broad range of applications. Using these models in a specific context typically relies on adjusting their context-specific parameters using our domain knowledge or some measurements/observations. For instance, finite element (FE) models can be used to simulate the behavior of many materials under a wide range of loading conditions [94]. However, for a specific application such as modeling the thermoforming process of woven carbon fiber reinforced plastics [95, 96, 28], a specific FE model is needed. The material parameters such as directional Young’s moduli and yarn-yarn friction coefficients in this FE model should be calibrated such that it can reproduce force-displacement curves obtained via experiments such as tension or three-point bending tests. Since such experimental data does not include the material parameters themselves, one must solve an inverse problem where the FE model’s calibration parameters are estimated such that the model fits the experimental data. During this process, it is implicitly presumed that the response (e.g., force-displacement curve) is sufficiently sensitive to the calibration parameters (e.g., Young’s moduli and yarn-yarn friction coefficients) as otherwise they cannot be accurately estimated.
Calibration of a computer model is tightly connected to that model’s bias with respect to an HF data source (e.g., experiments). To explain this connection, we note that most computer models suffer from systematic errors that arise from, e.g., their missing physics, the simplifying assumptions made during their development, or numerical errors. To mitigate the effect of these errors, computer models sometimes include a few additional calibration or tuning parameters that may not even correspond to any physical properties of the system. One example is the artificial viscosity parameter that is used to stabilize explicit solution methods that are needed when modeling dynamic processes such as fracture via the FE method. A related example is the calibration of physics-based reduced-order models (ROMs) [97, 98, 99, 27, 100, 101] that simplify expensive computer models (such as direct numerical simulations or DNS) to gain computational speedups. Such simplifications introduce some bias into the ROMs whose effects are typically mitigated by calibrating material parameters such that a ROM can reproduce small HF data obtained from DNS. That is, even if the material parameters are known, one may have to adjust them for ROMs.
@@
  • Simultaneous calibration of multiple ($>1$) computer models: We presume that the calibration parameters of these models correspond to some unobserved characteristics of a system. We make this assumptions since it is not optimal to jointly calibrate the tuning parameters of different models that are added to them for reasons besides characterizing unobserved features of a system (note also that the number of these tuning parameters generally varies across different models that simulate the same system).
@@
Following the notation of previous sections, we denote the quantitative inputs by $\boldsymbol{x}$ and the latent representations of qualitative inputs and the categorical source indicator variable by $\boldsymbol{h}$ and $\boldsymbol{z}$, respectively. These inputs are shared across all the data sources but, as described above, the LF sources have additional quantitative inputs that correspond to the calibration parameters and are denoted by $\boldsymbol{\zeta}=[\zeta_{1},\zeta_{2},\ldots,\zeta_{d_{\zeta}}]^{T}$. The “best” calibration parameters ($\boldsymbol{\zeta}^{*}$) should be estimated using the HF data to minimize the discrepancy between the HF and LF sources. We denote these estimates by $\widehat{\boldsymbol{\zeta}}$ and modify the correlation function to obtain them. For instance, the Gaussian and Matern correlation functions are reformulated as follows:
@@
where $\boldsymbol{u}=\mathopen{}\left[\boldsymbol{x},\boldsymbol{t},s,\boldsymbol{\zeta}\right]\mathclose{}^{T}$ and $\boldsymbol{\omega}$, $\boldsymbol{\theta}_{h}$ and $\boldsymbol{\theta}_{z}$ are defined as before. While training the model, the correlation between LF samples can be readily calculated via Equation 34b. However, if at least one of the samples is an HF one, in the last term of Equations 34a and 34b we use $\widehat{\zeta_{i}}$ which are estimated jointly with all the other parameters of the model via MAP.
Similar to Section 4.2, we can inversely learn the calibration parameters within a probabilistic setting to more accurately quantify the uncertainties compared to the deterministic counterpart based on the MAP. Following Section 4.2, we learn the calibration parameters within a variational framework by formulating their joint posterior with a multivariate normal distribution that is fully characterized by its mean vector and covariance matrix. We learn the parameters of this joint distribution via the reparametrization trick:
@@
where $\boldsymbol{\varepsilon}_{\zeta i}\sim\mathcal{N}(0,1)$ is an auxiliary noise variable while $\mu_{\zeta i}$ and $\tau_{\zeta i}$ parameterize the posterior distribution of $\zeta i$. During both the training and prediction phases, we draw samples for calibration parameters. As a result, the values of $\boldsymbol{\zeta}$ fluctuate in each optimization iteration which consequently changes the covariance and mean functions of the emulator. To efficiently consider these variations, we follow our method used for probabilistic manifold modeling (see Section 4.2) and employ ensembling to calculate both the mean vector and covariance matrix.
@@ 5 Functionalities of GP+  and Comparative Studies
Throughout this section, we use normalized root mean squared error (NRMSE) and normalized interval score (NIS) for assessing the accuracy of, respectively, the mean values and prediction intervals provided by a GP:
@@
where ${y}^{(i)}=y({\boldsymbol{u}}^{(i)})$ and ${\mu}^{(i)}=\mu{({\boldsymbol{u}}^{(i)})}$ denote the output and predicted mean of test sample ${\boldsymbol{u}}^{(i)}$, respectively, and $std(\boldsymbol{y})=std(y(\boldsymbol{U}))$ shows the standard deviation of the test samples obtained at $\boldsymbol{U}$. $\mathcal{U}^{(i)}$ and $\mathcal{L}^{(i)}$ are upper and lower endpoints of the prediction interval for the $i^{th}$ test sample. These endpoints are the predictive quantiles at levels $v/2$ and $21-v/$ 2, respectively. We use $95\%$ prediction interval ($v=0.05$) and hence these endpoints are defined as $\mathcal{U}^{(i)}={\mu}^{(i)}+1.96{\tau}^{(i)}$ and $\mathcal{L}^{(i)}={\mu}^{(i)}-1.96{\tau}^{(i)}$. $\mathbbm{1}\{\cdot\}$ is an indicator function which is $1$ if its condition holds and zero otherwise. For both metrics in Equations 36 and 37 lower values indicate more accuracy.
Unless otherwise stated, we use the default settings of GP+ which are detailed in Tables 9 and 10. For instance, the prior distributions of the parameters in all the examples are $\omega_{i}\sim N(-3,3)$, $\beta\sim N(0,1)$, $A_{ij}\sim N(0,1)$, $\sigma^{2}\sim LN(0,1)$ [^13] Log-Normal , and $\delta_{i}\sim LHS(0,0.01)$ [^14] Log-Half-Horseshoe with zero lower bound and scale parameter $0.01$. [102].
@@ 5.1 Emulation
We use three analytic and two engineering problems to compare the performance of GP+ against the GP emulation capabilities of GPytorch [^15] GPyTorch uses the Adam optimizer by default and to improve its performance we use a learning rate scheduler during the training process. To avoid convergence to local optima, we repeat the optimization process $64$ times, each with a different initialization for the model’s parameters., MATLAB [^16] We use automatic relevance determination (ARD) squared exponential kernel for GP emulation in MATLAB. and BoTorch [103, 104, 30]. These problems are briefly described below (see details in Appendix A) and they cover a range of input dimensionality and characteristics (e.g., HOIP only has categorical inputs).
As shown in Figure 8, emulation via GP+ is achieved via a few lines of code regardless of whether the problem has categorical variables or not. As opposed to GP+, MATLAB handles categorical variables by first one-hot encoding them and then treating the resultant variables as numerical. BoTorch leverages mixed single-task GP (MST-GP) which defines two distinct kernels for numerical and categorical features. Specifically, MST-GP uses the Matèrn kernel for the numerical features while for the categorical features it calculates the exponential of their normalized binary distance which is 00 when the two categorical variables are the same. The final kernel of MST-GP is the combination of the categorical and numerical parts (see Appendix D for details). Since MST-GP is specifically developed to handle mixed input spaces that have both categorical and numerical features, we use GPytorch in problems with only quantitative features (note that Gpytorch cannot handle categorical inputs).
@@
Wing and Borehole are two single-response analytic examples whose input space only has quantitative features. The dimensionality of the input space for Wing and Borehole is $10$ and $8$, respectively. To compare the performance of the three methods in mixed input spaces, we convert the first and sixth features of Borehole to categorical variables with $5$ distinct levels. This analytic example is referred to as Borehole-Mixed (see Section A.2 for further details). We generate $10000$ samples from the HF source in Table 4 and use $1\%$ of the data for training and the rest for testing. HOIP and NTA are both $3$-dimensional problems and only have categorical inputs (see Section A.1 for more details). We use $150$ and $100$ HF samples from HOIP and NTA, respectively, for emulation and the rest of the HF data for testing. For all problems, we repeat the emulation process $10$ times and report the average values for each metric to ensure the metrics are robust to random initialization.
Emulation results are summarized in Table 1 which demonstrate that GP+ more accurately predicts the responses in all examples and provides more reliable prediction intervals. In the case of Wing or Borehole which only have quantitative features, we attribute the superiority of GP+ primarily to the parameter optimization process. Specifically, we use MAP (as opposed to maximum likelihood estimation or MAE), search for the length-scale parameters (i.e., $\boldsymbol{\omega}$ in the correlation function) in the log scale, and leverage L-BFGS-B [^17] Limited-memory Broyden–Fletcher–Goldfarb–Shanno that considers simple bounds on the variables. for optimization. These choices smooth the profile of the objective function and accelerate the convergence.
@@
We observe a similar trend for Borehole-Mixed, HOIP, and NTA  which have categorical variables. In these problems, GP+ explicitly learns the relations between different categorical variables and their levels which not only improve the emulation performance, but also provide visually interpretable embeddings (see Figures 15(b), 15(a) and 15(c) in Section 5.4 for an example). We note GP+ estimate more parameters than other methods in Table 1 since it directly learns the correlations among categorical variables. This approach results in a more expensive and challenging optimization process which can converge to suboptimal solutions if the training data is very small and the categorical variables have many levels. To mitigate potential overfitting issues in such cases, we recommend using tighter priors in $f_{h}(\boldsymbol{\pi_{\boldsymbol{t}}},\boldsymbol{\theta}_{h})$. Correspondingly, in NTA and HOIP with $240$ and $480$ distinct categorical combinations, respectively, we use $\mathcal{N}(0,0.1)$ and $\mathcal{N}(0,0.01)$ priors for $f_{h}(\boldsymbol{\pi_{\boldsymbol{t}}},\boldsymbol{\theta}_{h})$.

### v2

@@ 4.3 Gaussian Processes with Mixed Basis Functions
The parametric mean and covariance functions in Equation 3b can be formulated in many ways. In this regard, most advancements have focused on designing the kernel (e.g., the ones we develop in Equations 13b and 25 or deep kernels [lee2017deep, al2017learning]) since it significantly affects the performance of the resulting GP model. However, the mean function in Equation 3a plays an important role in many applications that involve, e.g., extrapolation [planas2021evolutionary], fusing multi-source data, or identifying model form errors.
Existing techniques typically leverage polynomials (in combination with other analytic functions such as $\sin(\cdot),\log(\cdot),\cdots$) or NNs in designing $m(\boldsymbol{x};\boldsymbol{\beta})$. In GP+, we extend these methods to seamlessly include the categorical variables in the mean function. Specifically, our idea is to feed the learnt representations of $\boldsymbol{t}$ and $s$ into the mean function instead of the original categorical variables, i.e., we reparameterize $m(\boldsymbol{x},\boldsymbol{t},s;\boldsymbol{\beta})$ to $m(\boldsymbol{x},\boldsymbol{h},\boldsymbol{z};\boldsymbol{\beta})$, see Figure 5. A major difference between our reparameterization and other alternatives (such as an NN whose inputs are one-hot encoded representation of $\boldsymbol{t}$ and $s$) is that our mean and covariance functions are directly coupled since the latent variables used in $m(\boldsymbol{x},\boldsymbol{h},\boldsymbol{z};\boldsymbol{\beta})$ are parameterized in the kernel, i.e., $\boldsymbol{h}=f_{h}(\boldsymbol{\pi}_{t};\boldsymbol{\theta}_{h})$ and $\boldsymbol{z}=f_{z}(\boldsymbol{\pi}_{s};\boldsymbol{\theta}_{z})$. Based on this idea, in GP+ we provide the following two options for modeling the mean function: (1) Having a global function that is shablack among all combinations of $\boldsymbol{t}$ and $s$, and (2) Having mixed basis functions where a unique mean function is learnt for specific combinations of the categorical variables (e.g., in MF modeling, we can learn a unique mean function for each of the $s$ data sources).
Figure 5: Multi-fidelity modeling via mixed basis functions: Two generic options are defined in GP+ for building the mixed bases: $(1)$ pblacketermined bases where multiple bases like polynomial, $sin(\cdot)$ and $cos(\cdot)$ can be defined for each data source, $(2)$ FFNNs with user-defined architectures. All the parameters of the mean functions have a normal prior and are jointly learned through MAP.
Figure 5 illustrates two generic options that we define in GP+ for building mixed bases. The first option builds the mean function based on pre-determined bases that can include polynomials, $sin(\cdot),log(\cdot)$ or any other analytic functions. The second option is based on a fully connected FFNN whose architecture (e.g., number of hidden layers and their sizes) should be designed by the user. The size of the input layer of the NN depends on the dimensionality of $\boldsymbol{x},\boldsymbol{h},$ and $\boldsymbol{z}$ while its output layer size depends on the response dimensionality (hence the output size is $1$ for a single-response dataset).
@@
The results are illustrated in Figure 6 and indicate that the second GP emulates both sources better than the first GP in both interpolation and extrapolation. As it can be seen in Table 4 the true model form error is $10.3x^{2}-0.7x+$ 1 while the discoveblack one with the second GP is $0.99390.2981x^{2}-0.7059x+$ 0.9939. We attribute the small differences between these two functions primarily to the fact that the training data is very small and noisy. As also shown in Figure 6 we observe that the inclusion of the mixed bases affects the learnt encoding for $s$ whose two levels are mapped to distant latent points in the first GP but close-by points in the second GP. This behavior indicates that the entire model form error in the second GP can be obtained by comparing the mean functions associated with the two sources.
@@ 4.4 Inverse Parameter Learning for Computer Models
Most computer models are built to be applicable to a broad range of applications. Using these models in a specific context typically relies on adjusting their context-specific parameters using our domain knowledge or some measurements/observations. For instance, finite element (FE) models can be used to simulate the behavior of many materials under a wide range of loading conditions [RN311]. However, for a specific application such as modeling the thermoforming process of woven carbon fiber reinforced plastics [10.1063/1.4963592, BOTELHO20031843, RN791], a specific FE model is needed. The material parameters such as directional Young’s moduli and yarn-yarn friction coefficients in this FE model should be calibrated such that it can reproduce force-displacement curves obtained via experiments such as tension or three-point bending tests. Since such experimental data does not include the material parameters themselves, one must solve an inverse problem where the FE model’s calibration parameters are estimated such that the model fits the experimental data. During this process, it is implicitly presumed that the response (e.g., force-displacement curve) is sufficiently sensitive to the calibration parameters (e.g., Young’s moduli and yarn-yarn friction coefficients) as otherwise they cannot be accurately estimated.
Calibration of a computer model is tightly connected to that model’s bias with respect to an HF data source (e.g., experiments). To explain this connection, we note that most computer models suffer from systematic errors that arise from, e.g., their missing physics, the simplifying assumptions made during their development, or numerical errors. To mitigate the effect of these errors, computer models sometimes include a few additional calibration or tuning parameters that may not even correspond to any physical properties of the system. One example is the artificial viscosity parameter that is used to stabilize explicit solution methods that are needed when modeling dynamic processes such as fracture via the FE method. A related example is the calibration of physics-based blackuced-order models (ROMs) [RN1182, RN1655, RN1854, RN1786, dvorak1992transformation, roussette2009nonuniform] that simplify expensive computer models (such as direct numerical simulations or DNS) to gain computational speedups. Such simplifications introduce some bias into the ROMs whose effects are typically mitigated by calibrating material parameters such that a ROM can reproduce small HF data obtained from DNS. That is, even if the material parameters are known, one may have to adjust them for ROMs.
@@
  • Simultaneous calibration of multiple ($>1$) computer models: We presume that the calibration parameters of these models correspond to some unobserved characteristics of a system. We make this assumption since it is not optimal to jointly calibrate the tuning parameters of different models that are added to them for reasons besides characterizing unobserved features of a system (note also that the number of these tuning parameters generally varies across different models that simulate the same system).
@@
Following the notation of previous sections, we denote the quantitative inputs by $\boldsymbol{x}$ and the latent representations of qualitative inputs and the categorical source indicator variable by $\boldsymbol{h}$ and $\boldsymbol{z}$, respectively. These inputs are shablack across all the data sources but, as described above, the LF sources have additional quantitative inputs that correspond to the calibration parameters and are denoted by $\boldsymbol{\zeta}=[\zeta_{1},\zeta_{2},\ldots,\zeta_{d_{\zeta}}]^{T}$. The “best” calibration parameters ($\boldsymbol{\zeta}^{*}$) should be estimated using the HF data to accurately characterize the physical system. We denote these estimates by $\widehat{\boldsymbol{\zeta}}$ and modify the correlation function to obtain them. For instance, the Gaussian and Matérn correlation functions are reformulated as follows:
@@
where $\boldsymbol{u}=\mathopen{}\left[\boldsymbol{x},\boldsymbol{t},s,\boldsymbol{\zeta}\right]\mathclose{}^{T}$ and $\boldsymbol{\omega}$, $\boldsymbol{\theta}_{h}$ and $\boldsymbol{\theta}_{z}$ are defined as before. While training the model, the correlation between LF samples can be readily calculated via Equation 28b. However, if at least one of the samples is an HF one, in the last term of Equations 28a and 28b we use $\widehat{\zeta_{i}}$ which are estimated jointly with all the other parameters of the model via MAP.
Similar to Section 4.2, we can inversely learn the calibration parameters within a probabilistic setting to more accurately quantify the uncertainties compablack to the deterministic counterpart based on the MAP. Following Section 4.2, we learn the calibration parameters within a variational framework by formulating their joint posterior with a multivariate normal distribution that is fully characterized by its mean vector and covariance matrix. We learn the parameters of this joint distribution via the reparametrization trick:
@@
where $\boldsymbol{\varepsilon}_{\zeta i}\sim\mathcal{N}(0,1)$ is an auxiliary noise variable while $\mu_{\zeta i}$ and $\tau_{\zeta i}$ parameterize the posterior distribution of $\zeta i$. During both the training and pblackiction phases, we draw samples for calibration parameters. As a result, the values of $\boldsymbol{\zeta}$ fluctuate in each optimization iteration which consequently changes the covariance and mean functions of the emulator. To efficiently consider these variations, we follow our method used for probabilistic manifold modeling (see Section 4.2) and employ ensembling to calculate both the mean vector and covariance matrix.
We highlight that in both deterministic and probabilistic calibration cases, the estimated calibration parameters and the learnt bias are tightly connected in that the former depends on what bias form has been chosen. Most existing methods [RN377, RN375, RN271, RN272, RN335, RN336, RN645, RN607, RN295, RN705] first assume a specific functional form (e.g., a GP or a polynomial) for the bias and the relation between the LF and HF sources (see Equation 14 and Figure 3 for one example). Then, given data from both LF and HF sources, they estimate the calibration parameters and the parameters of the bias function. For these approaches, the estimated calibration parameters are strongly dependent on the assumed form of the bias function and how it relates the HF and LF sources. If these assumptions are incorrect, the calibration results will be misleading. In GP+ we significantly relax these assumptions. Specifically, we $(1)$ do not assume the bias term is additive, and $(2)$ simultaneously calibrate multiple sources (rather than just calibrating one source at a time). Therefore, we do not eliminate the so-called identifiability issue but provide the means that analysts can use to address it depending on the application. For instance, an effective way to blackuce identifiability issues is using multiple-response data during calibration [RN607, RN272, RN271, RN277]. Similarly, the multi-source calibration mechanism in GP+ provides the calibration process with more information and hence has the potential to blackuce non-identifiability. Additionally, in GP+ we can use mixed basis functions which can help analysts in choosing appropriate mean functions for each source and study the effects of this choice on the estimated calibration parameters, accuracy on unseen data, and learnt bias functions (note that the difference between two mean functions essentially gives the global bias between the corresponding sources, see Figure 6).
@@ 5 Functionalities of GP+ and Comparative Studies
Throughout this section, we use normalized root mean squablack error (NRMSE) and normalized interval score (NIS) for assessing the accuracy of, respectively, the mean values and pblackiction intervals provided by a GP:
@@
where ${y}^{(i)}=y({\boldsymbol{u}}^{(i)})$ and ${\mu}^{(i)}=\mu{({\boldsymbol{u}}^{(i)})}$ denote the output and pblackicted mean of test sample ${\boldsymbol{u}}^{(i)}$, respectively, and $std(\boldsymbol{y})=std(y(\boldsymbol{U}))$ shows the standard deviation of the test samples obtained at $\boldsymbol{U}$. $\mathcal{U}^{(i)}$ and $\mathcal{L}^{(i)}$ are upper and lower endpoints of the pblackiction interval for the $i^{th}$ test sample. These endpoints are the pblackictive quantiles at levels $v/2$ and $21-v/$ 2, respectively. We use $95\%$ pblackiction interval ($v=0.05$) and hence these endpoints are defined as $\mathcal{U}^{(i)}={\mu}^{(i)}+1.96{\tau}^{(i)}$ and $\mathcal{L}^{(i)}={\mu}^{(i)}-1.96{\tau}^{(i)}$. $\mathbbm{1}\{\cdot\}$ is an indicator function which is $1$ if its condition holds and zero otherwise. For both metrics in Equations 30 and 31 lower values indicate more accuracy.
Unless otherwise stated, we use the default settings of GP+ which are detailed in Tables 9 and 10. For instance, the prior distributions of the parameters in all the examples are $\omega_{i}\sim N(-3,3)$, $\beta\sim N(0,1)$, $A_{ij}\sim N(0,1)$, $\sigma^{2}\sim LN(0,1)$ [^14] Log-Normal, and $\delta_{i}\sim LHS(0,0.01)$ [^15] Log-Half-Horseshoe with zero lower bound and scale parameter $0.01$. [carvalho2010horseshoe].
@@ 5.1 Emulation
We use three analytic and two engineering problems to compare the performance of GP+ against the GP emulation capabilities of GPytorch [^16] GPyTorch uses the Adam optimizer by default and to improve its performance we use a learning rate scheduler during the training process. To avoid convergence to local optima, we repeat the optimization process $64$ times, each with a different initialization for the model’s parameters., MATLAB [^17] We use automatic relevance determination (ARD) squablack exponential kernel for GP emulation in MATLAB., BoTorch, and SMT2 [Matlab, BoTorch, RN1270, saves2024smt]. These problems are briefly described below (see details in Appendix A) and they cover a range of input dimensionality and characteristics (e.g., HOIP only has categorical inputs). To provide a comprehensive analysis, we illustrate the emulation comparisons in this section and discuss the computational costs of these baselines in Appendix I.
[…]
@@
Wing and Borehole are two single-response analytic examples whose input space only has quantitative features. The dimensionality of the input space for Wing and Borehole is $10$ and $8$, respectively. To compare the performance of the three methods in mixed input spaces, we convert the first and sixth features of Borehole to categorical variables with $5$ distinct levels. This analytic example is referblack to as Borehole-Mixed (see Section A.2 for further details). We generate $10000$ samples from the HF source in Table 4 and use $1\%$ of the data for training and the rest for testing. HOIP and NTA are both $3$-dimensional problems and only have categorical inputs (see Section A.1 for more details). We use $150$ and $100$ HF samples from HOIP and NTA, respectively, for emulation and the rest of the HF data for testing. For all problems, we repeat the emulation process $10$ times and report the average values for each metric to ensure the metrics are robust to random initialization.
Emulation results are summarized in Table 1 which demonstrate that GP+ more accurately pblackicts the responses in all examples and provides more reliable pblackiction intervals. In the case of Wing or Borehole which only have quantitative features, we attribute the superiority of GP+ primarily to the parameter optimization process. Specifically, we use MAP (as opposed to maximum likelihood estimation or MAE), search for the length-scale parameters (i.e., $\boldsymbol{\omega}$ in the correlation function) in the log scale, and leverage L-BFGS-B [^18] Limited-memory Broyden–Fletcher–Goldfarb–Shanno that considers simple bounds on the variables.. for optimization. These choices smooth the profile of the objective function and accelerate the convergence. The comparable performance of SMT2 and GP+ in terms of NRMSE is attributed to SMT2’s usage of the so-called profiling technique [RN783] for parameter estimation.
@@
GP+ outperforms all baselines in Borehole-Mixed, HOIP, and NTA  which have categorical variables. In these problems, GP+ explicitly learns the relations between different categorical variables and their levels which not only improve the emulation performance, but also provide visually interpretable embeddings (see Figures 16(b), 16(a) and 16(c) in Section 5.4 for an example). We note GP+ estimate more parameters than other methods in Table 1 since it directly learns the correlations among categorical variables. This approach results in a more expensive and challenging optimization process which can converge to suboptimal solutions if the training data is very small and the categorical variables have many levels. To mitigate potential overfitting issues in such cases, we recommend using tighter priors in $f_{h}(\boldsymbol{\pi_{\boldsymbol{t}}},\boldsymbol{\theta}_{h})$. Correspondingly, in NTA and HOIP with $240$ and $480$ distinct categorical combinations, respectively, we use $\mathcal{N}(0,0.1)$ and $\mathcal{N}(0,0.01)$ priors for $f_{h}(\boldsymbol{\pi_{\boldsymbol{t}}},\boldsymbol{\theta}_{h})$.